OpenStudy (anonymous):

19 math questions. 7th grade. Need help!!

OpenStudy (anonymous):

Give us the question..?

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Which is a counterexample to this conjecture? The sum of any two consecutive integers is a composite number. A. 16 + 17 = 33 B. 10 + 11 = 21 C. 6 + 7 = 13 D. 7 + 8 = 15

jimthompson5910 (jim_thompson5910):

Which result is a prime number?

OpenStudy (anonymous):

idk:/

jimthompson5910 (jim_thompson5910):

a prime number is a number that only has factors of 1 and itself example: 33 is not prime because 3*11 = 33, but 37 is prime because only 1 and 37 are factors

OpenStudy (anonymous):

Im confuzzled.

jimthompson5910 (jim_thompson5910):

this might help https://www.mathsisfun.com/prime_numbers.html

OpenStudy (anonymous):

Okay. Thats prime numbers. What about this problem?

jimthompson5910 (jim_thompson5910):

Well a number is either composite or it is prime. It can't be both. If it's not composite, then it is prime

jimthompson5910 (jim_thompson5910):

The claim is `The sum of any two consecutive integers is a composite number.` but there is one case in your answer choices where 2 consecutive integers add to a prime number

OpenStudy (anonymous):

This problem still dont make any sense.

jimthompson5910 (jim_thompson5910):

A composite number is any number that has factors other than 1 and itself example: 36 is composite because 3*12 = 36

jimthompson5910 (jim_thompson5910):

do you see in your answer choices 2 numbers being added to get a prime number?

OpenStudy (anonymous):

Yes. C.

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

sorry my notifications aren't working

jimthompson5910 (jim_thompson5910):

Yes `C. 6 + 7 = 13` is where you add 2 consecutive integers and you get a number that is NOT composite So the claim `The sum of any two consecutive integers is a composite number.` is proven false with choice C being a counter example